cost R353500 , overall cost R45000 and mark up 22% on sale

how much profit?
1
Expert's answer
2013-02-12T09:28:55-0500
CostR35350, overall cost R45000 and mark up 22% sale means that cost with sale will
be R45000-R45000*22%=R45000-R45000*0,22=R45000-R9900=R35100. It means that with
sale it is cheaper than without sale. Profit is R35350-R35100=R250.
Answer: R250.